OpenStudy (anonymous):

y=2/5x+3/15 both divided by -1/5

OpenStudy (anonymous):

I'm trying to et the equation solved so I can graph it.

OpenStudy (texaschic101):

well.....2/5 divided by -1/5 = 2/5 * -5/1 = - 10/5 = -2 (slope) 3/15 divided by -1/5 = 3/15 * -5/1 = - 15/15 = -1 (y int)

OpenStudy (anonymous):

Woah.

OpenStudy (texaschic101):

so y = -2x - 1 y int is -1.....(0,-1) x int can be found by subbing in 0 for y 0 = -2x - 1 2x = -1 x = -1/2.....(-1/2,0) graph : C

OpenStudy (anonymous):

Thanks.

OpenStudy (texaschic101):

YW :)

OpenStudy (anonymous):

Can you help me with 2 more problems?

OpenStudy (texaschic101):

I am sorry but I am gonna have to go now...I kinda lost track of time

OpenStudy (texaschic101):

I will be back on later

OpenStudy (texaschic101):

so so so so sorry

OpenStudy (anonymous):

Lol its ok. But thanks. :)